OpenStudy (anonymous):

In which pair of points does the graph of y=2/3x-5 pass through? a. (3,-3) , (-6,-9) b. (-3,-3) , (6,9) c. (0,3) , (6,9) d. (3,6) , (9,0)

OpenStudy (anonymous):

same method try each pair just plug the points in the equation the pair which will satisfy is the answer for example lets try pair a ) (3,-3) , (-6,-9) let me simplify equation a bit 5+y= 2/3x now put x=3 and y =-3 5 +(-3) =2/3(3) 2=2/9 or 2*9=2 18=2 which is incorrect (you can try the second point but it will not be answer as one of the point is incorrect ) adopt similar procedure for options b, c and d can you solve this now ?

OpenStudy (anonymous):

i'll try :D

OpenStudy (anonymous):

how can you substitute ? there are 2 points?

OpenStudy (anonymous):

wait !! i understand now xD hahah is the answer letter a?

OpenStudy (anonymous):

you substitite turn by turn for example inorder to test option b there re two points (-3,-3) , (6,9) first test (-3,-3) equation (simplified) is 5+y= 2/3x plug in the point 5-3=2/3(-3) 2=-2/9 2*9=-2 18= -2 which is not true now test the second point (6,9) put in the equation 5+y= 2/3x 5+9=2/3(3) 14=2/9 which is again incorrect . so move to option C

OpenStudy (anonymous):

wait did you write the equation correctly ? is it \[\Large y=\frac{2}{3x}-5\] or \[\Large y=\frac{2}{3x-5}\] which one ?

OpenStudy (anonymous):

the 1st one

OpenStudy (anonymous):

ok its ok . exactly what I interpreted at first . Try option c and d

OpenStudy (anonymous):

Sorry a is not correct give it another shot ! :)

OpenStudy (anonymous):

okay im going to answer c and d

OpenStudy (anonymous):

i didn't get the right answers

OpenStudy (anonymous):

same here I don't see the points which satisfy the equation

OpenStudy (anonymous):

are you sure the equation is \[\Large y=\frac{2}{3x}-5\] ???????

OpenStudy (anonymous):

\[y=\frac{ 2 }{ 3 }x-5\]

OpenStudy (anonymous):

oppss

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

a is the right answer right?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

haha xD sorry btw

OpenStudy (anonymous):

its ok I knew something is wrong with equation :)